A sports car must achieve a velocity of 100m/s after passing the entry tollgate of an expressway. if it can do this travelling 1 km, what must its acceleration be?
1
Expert's answer
2021-10-07T08:50:32-0400

"v_0=0"

"v_1 = 100\frac{m}{s}"

"s =1km =1000m"

"s= \frac{v_1^2-v_0^2}{2a}"

"a= \frac{v_1^2-v_0^2}{2s}=\frac{100^2-0^2}{2*1000}=5\frac{m}{s^2}"


"\text{Answer: }5\frac{m}{s^2}"